OpenStudy (anonymous):

solve the eponential equation: 5^x = 4^(x+1)

OpenStudy (anonymous):

5^x=4^(x+1) ln(5^x)=ln(4^x+1) xln(5)=(x+1)ln(4) x/(x+1)=ln(4)/ln(5) or you can sove it as x/x+1)=0.8614 and without rounding and choping x/(x+1)=0.861353116

OpenStudy (anonymous):

As above then So x=6.212567